# Project page: https://github.com/zhang-tao-whu/PCM

### v3

@@ Abstract
Recently, state space models have exhibited strong global modeling capabilities and linear computational complexity in contrast to transformers. This research focuses on applying such architecture in point cloud analysis. In particular, for the first time, we demonstrate that Mamba-based point cloud methods can outperform previous methods based on transformer or multi-layer perceptrons (MLPs). To enable Mamba to process 3-D point cloud data more effectively, we propose a novel Consistent Traverse Serialization method to convert point clouds into 1-D point sequences while ensuring that neighboring points in the sequence are also spatially adjacent. Consistent Traverse Serialization yields six variants by permuting the order of x, y, and z coordinates, and the synergistic use of these variants aids Mamba in comprehensively observing point cloud data. Furthermore, to assist Mamba in handling point sequences with different orders more effectively, we introduce point prompts to inform Mamba of the sequence’s arrangement rules. Finally, we propose positional encoding based on spatial coordinate mapping to inject positional information into point cloud sequences better. Point Cloud Mamba surpasses the state-of-the-art (SOTA) point-based method PointNeXt and achieves new SOTA performance on the ScanObjectNN, ModelNet40, ShapeNetPart, and S3DIS datasets. It is worth mentioning that when using a more powerful local feature extraction module, our PCM achieves 82.6 mIoU on S3DIS, significantly surpassing the previous SOTA models, DeLA and PTv3, by 8.5 mIoU and 7.9 mIoU, respectively. Code and model are available at https://github.com/SkyworkAI/PointCloudMamba.
Keywords:
3D Point Cloud State Space Models Visual Mamba
@@
Figure 1: Several pipelines of point cloud modeling. (a) denotes point-based methods with only local perception, including point-based methods, such as PointNet [38], PointNet++ [39], PointMLP [35], and PointNeXt [41]. (b) is the transformer-based method with global perception but quadratic computational cost, including Point Transformer [75] and Point-MAE [37]. (c) represents Mamba-based methods, which offer advantages of global modeling and linear computational complexity.
1 Introduction
@@
Point cloud analysis [38, 39, 27, 75, 41, 58, 41] has become a popular topic in 3D understanding and has drawn attention from the research community. Unlike 2D image processing, point clouds are composed of unordered and irregular point sets, making it difficult to apply the 2D image processing methods directly. Thus, recent deep-learning-based approaches [38, 39, 27, 58, 35, 41, 75, 15] propose using various methods, such as voxel-based and point-based, for point cloud representation. The representative works, Point-Net [38] and Point-Net++ [39], adopt the MLP-based design with deep hierarchical local priors, as shown in Fig. 1 (a). After that, many research works [35, 3, 48, 54, 75, 15, 60, 59] focus on advanced local geometric modeling via convolution, graph modeling, or attention.
@@
Meanwhile, with the rapid progress of vision transformers, several works [75, 15, 60, 59] enhance the global modeling with transformer structure in point could, as shown in Fig. 1 (b). In addition, transformer architectures also work effectively in mask point pre-training, 3D segmentation, and in-context learning. However, the computation and memory costs are still huge. Recently, state space models [13, 14] (SSMs) have been proven to model long-range dependency in sequential data. In particular, Mamba [13] is proven effective as Transformer [53] for several challenging NLP tasks. After that, recent works [77, 31, 25, 2, 34, 66, 70, 43, 16] explore SSMs in various vision tasks, including image representation learning, medical segmentation, and low-level vision tasks. One concurrent work, PointMamba, uses the Mamba layer to model the global context. However, there are still significant performance gaps between PointMamba and previous point-based methods.
@@
In this work, we ask an essential question: Can we design an efficient point cloud analysis architecture using Mamba and surpass the performance of point-based methods? In particular, we introduce the Point Cloud Mamba (PCM), a combining local and global modeling framework that outperforms the SOTA point-based method PointNeXt [41].
@@
PCM utilizes Mamba architecture to model the global features of point clouds while maintaining linear computational complexity. To process 3-D point cloud data effectively by Mamba layers, we propose a novel Consistent Traverse Serialization (CTS) method to serialize point clouds into a 1-D point sequence while ensuring that neighboring points in the sequence are also adjacent in space. Then, CTS can easily derive six variants by simply permuting the order of x, y, and z coordinates. Additionally, when these six variants of CTS are combined, Mamba layers can more effectively model point cloud features. This is because the different variants provide various perspectives of the point cloud, ensuring that spatially adjacent points are also adjacent in a serialized point sequence. To help Mamba handle specific point sequences better, we introduce order prompts to provide Mamba with the arrangement rules of the current point sequence. Finally, we propose simple spatial coordinate mapping as positional embedding for points, more suitable for irregular point cloud data than RoPE [49] and learnable embedding.
@@
Thanks to the above improvements, we successfully introduced Mamba into Point Cloud analysis and obtained Point Cloud Mamba (PCM). PCM outperforms the SOTA point-based method PointNeXt on three datasets: ScanObjectNet [52], ModelNet40 [63], and ShapeNetPart [72]. On the ShapeNetPart part segmentation dataset, PCM-Tiny surpasses PointNeXt with a 0.4 OA improvement, using only 39% of the parameter of PointNeXt. Using the more powerful local feature extractor DeLA [6], PCM achieved 82.6 mIoU on the S3DIS dataset, significantly surpassing PTv3 [59] by 7.9 mIoU.
@@
In summary, we have the following contributions: 1) We introduce Mamba into point cloud analysis and construct a combined local and global modeling framework named Point Cloud Mamba. 2) We propose consistent traverse serialization, order prompts, and positional encoding based on spatial coordinate mapping to assist Mamba in better handling point cloud data. 3) Point Cloud Mamba is the first Mamba-based method that works well in point cloud analysis. It outperforms the SOTA point-based method PointNeXt on ScanObjectNet, ModelNet40, ShapeNetPart, and S3DIS datasets.
2 Related Work
@@
3D Point Cloud Classification. Recent works have used deep neural networks to process 3D point clouds. In particular, representative works, PointNet [38] and PointNet++ [39], are the pioneering point-based approaches to handle the point clouds using MLPs directly. Meanwhile, several works explore graph-based modeling to utilize 3D geometric topology. Then, several works [58, 51, 47, 27, 67, 21, 12, 8, 78] explore the local geometric features via different kernel modeling. Moreover, several works [30, 68, 64, 42, 7, 19, 29, 65, 74, 35, 24, 9, 69] explore other point cloud architecture designs, including MLPs and transformers. Several works also explore different pre-training methods [37, 73, 44, 45, 57, 18, 20, 62, 76, 71] or in-context abilities [11, 56, 61] inspired by the NLP field. Recently, state space models [14, 13] have achieved significant progress. Compared with transformers, they have advantages in efficient global modeling. A concurrent work [28] explores such architecture in point clouds. However, there are still significant performance gaps compared with previous point cloud methods. Our PCM shows that Mamba architecture can achieve comparable or even better results than transformer-based models in 3D point clouds.
@@
3D Visual Transformers. With the rise of the transformer in 2D version [10, 26, 4], several works [22, 15, 46, 50, 23, 32, 55] also explore transformer architectures in the point cloud. Earlier works [15, 75] have focused on the point cloud process. PCT [15] performs global attention directly to each point, following the ViT [10]. However, it has memory consumption and computational complexity issues. Point Transformer [75] solves this issue by introducing local attention. Then, the updated versions [60, 59] explore the different architectures to improve performance and efficiency. Inspired by these studies, our works combine local point processing and a new traverse serialization strategy, which leads to better results than direct SSM traverse.
@@
State Space Models. Inspired by continuous state space models in control systems, recently, state space models [13, 14] have been proven to model long-range dependency. In particular, S4 [14] proposes to normalize the parameter into the diagonal structure, which results in less computation cost and memory usage. After that, Mamba [13] presents a selection mechanism that leads to better results than transformers. Recently, several works have explored such architecture in different tasks, including image classification [77, 31, 25], graph modeling [2], medical segmentation [34, 66, 70, 43], and low-level version tasks [16]. As a concurrent work, we further prove the potential of SSMs in the 3D point clouds, where we can achieve even better results than previous architectures.
3 Method
@@
The SSM-based architecture, Mamba [13], is attractive for point cloud representation learning due to its global modeling capability and linear computational complexity. However, Mamba is designed for the causal modeling of 1-D sequences, making it difficult to be directly applied to the modeling of non-causal 3-D point cloud data, posing many challenges to be addressed.
This section explores how to effectively integrate Mamba into architectures based on local modeling to capture global features. We first review PointMLP [35], a local modeling architecture, and Mamba, a global modeling architecture with linear complexity, in Sec. 3.1. Then, in Sec. 3.2, we introduce how to combine Mamba with PointMLP to obtain a point cloud network based on Mamba architecture. Finally, we introduce improving this naive Mamba-based network to Point Cloud Mamba in Sec. 3.3. We propose consistent traverse serialization, order prompt, new positional embedding, and more reasonable architectures to assist Mamba in better modeling point cloud data.
3.1 Preliminaries
PointMLP formulation. PointMLP [35] models the representation of point clouds using simple MLP and expands the receptive field of each point through point cloud downsampling and local feature aggregation. The process can be described by Equ. 1 and 2:
@@ 3.1 Preliminaries
$GAM(\{f_{i,j}^{l}\})=\alpha\odot\frac{\{f_{i,j}^{l}\}-f_{i}^{l}}{\sigma+\delta},\quad\sigma=\sqrt{\frac{1}{k\times n\times d}\sum_{i=1}^{n}\sum_{j=1}^{k}(f_{i,j}^{l}-f_{i}^{l})^{2}},$ (2)
@@
Inspired by Equ. 4 and 5, Mamba [13], a new SSM-based model that introduces time-varying system parameters, has been introduced. Specifically, $\Delta$, $\overline{\mathbf{A}}$, and $\overline{\mathbf{B}}$ are all functions of $x_{t}$. Mamba demonstrates long sequence modeling capabilities comparable to Transformer and achieves linear computational complexity during inference following Equ. 5.
However, Equ. 5 is difficult to compute in parallel. It can be expanded and implemented using global convolution to enhance the efficiency of training Mamba on GPUs:
@@
3.2 A Naive Mamba-based Point Cloud Network
Architecture. Using Mamba [13] to replace the MLP operator in PointMLP [35] to achieve global modeling capabilities is a natural idea. However, this still requires addressing two challenges: 1) Point clouds are 3-D data, so how can we transform them into 1-D sequences? 2) Mamba is designed for causal modeling, so how can Mamba handle non-causal point cloud data? To address these challenges, we first adopt the z-order [36] serialization method to flatten 3-D point cloud data into 1-D sequences, allowing point cloud data to be processed by Mamba. Secondly, inspired by [77] and [31], we use bidirectional Mamba to allow each point to obtain features from any other point. At this point, we have implemented a naive Mamba-based network for the point cloud:
$f_{i}^{l+1}=Mamba(x^{l},reverse(x^{l})),\quad x^{l}=\mathcal{S}_{z}(\mathcal{A}(GAM(\{f_{i,j}^{l}|j=1,...,k\})))$ (7)
@@ 3.2 A Naive Mamba-based Point Cloud Network
Shortcomings. However, the naive Mamba-based network still has some shortcomings, only achieving 84.1 OA on ScanObjectNN, significantly underperforming modern point-based methods like PointNeXt. There are some reasons: 1) Using a single serialization method to convert 3-D point cloud data into 1-D point sequences significantly loses the spatial relational information between points. Specifically, even if a point is adjacent to many others in space, it can only be adjacent to adjacent points in the point sequence. Therefore, employing multiple serialization methods in combination to alleviate this loss of spatial relational information is necessary. 2) The architecture design of the naive Mamba-based network follows the point-based method PointMLP, which may not be optimal or even reasonable for Mamba-based architecture. Exploring what benefits Mamba in modeling point clouds is important and necessary. Please refer to the Sec. 4.3 for a detailed discussion.
3.3 Point Cloud Mamba
In Sec. 3.2, we obtained a naive mamba-based point cloud network; however, there is still plenty of room for optimization. Next, we will elaborate on improving this naive architecture to Point Cloud Mamba (PCM) and achieving performance beyond PointNeXt [41].
@@ 3.3 Point Cloud Mamba
Serialization strategy. How to convert 3-D point cloud data into a 1-D sequence that Mamba can handle is crucial. We find that Mamba can better process point cloud sequences arranged according to specific rules than disordered point cloud sequences. For example, randomly flattening point cloud data into a 1-D point sequence and feeding it into Mamba for modeling will significantly underperform compared to using an ordered point sequence (84.1 vs. 86.7). Based on this, we propose the Consistent Traverse Serialization (CTS) strategy, which ensures that adjacent points in the sequence are also adjacent in spatial position. Fig. 3 illustrates how CTS converts 3-D point clouds into a 1-D sequence.
@@
$Code\_func(n_{1},n_{2})=\left\{\begin{aligned} n_{2}\times N+n_{1},\quad\quad&n_{2}\%2=0\\ (n_{2}+1)\times N-n_{1},\quad&n_{2}\%2\neq 0\\ \end{aligned}\right.$ (9)
@@
Sorting the point cloud according to the code allows it to be flattened into a 1-D sequence. This simple serialization strategy performs comparably to carefully designed z-order [36] and Hilbert-order [17] serialization strategies. Additionally, by exchanging the order of the x, y, and z axes, six different serialization methods can be derived, which we call "xyz", "xzy", "yxz", "yzx", "zxy", and "zyx". These different serialization methods can be viewed as various point cloud observations from different spatial perspectives.
Additionally, combining multiple serialization strategies can effectively assist Mamba in better modeling point cloud features, as shown in Tab. 10. Specifically, we adopt different serialization strategies for the inputs of different Mamba layers, allowing Mamba to perceive the point cloud more comprehensively, thus significantly surpassing a single serialization strategy alone.
Figure 4: The order prompts. Different colors represent different serialization orders. $N_{p}$ order prompts are mapped to the same channel size as the features and then concatenated to the beginning and end of the point sequence input to the mamba layer.
@@
As shown in Fig. 4, we assigned $N_{p}$ learnable embeddings as order prompts for each serialization order. Before being processed by the Mamba layer, the point cloud sequence has the corresponding order prompts added to both the beginning and end of the sequence. Considering that the input feature dimensions of the Mamba layers at different stages may vary, we also allocate a shared Linear layer for all Mamba layers within each stage to map the order prompts to the required channel size.
Positional embedding. In sequence modeling, positional encoding is crucial and widely applied in text and image patch sequence modeling. For example, RoPE [49] and learnable positional embedding are widely used in text sequence modeling and image patch sequence modeling, respectively. However, we find that these positional encoding methods are unsuitable for point clouds due to their sparsity and irregular shapes. The spatial distances between any two adjacent points in a point cloud sequence vary significantly, making it difficult to use the sequence positional information to represent the spatial gaps between adjacent points accurately. We employ a shared positional mapping function to address this challenge. It maps the point spatial coordinates to positional embeddings with the same channel size as the features.
$Emd_{pos}=Linear(\{c_{1}^{s},c_{2}^{s},c_{3}^{s}\})\vspace{-2mm}$ (11)
@@
Architecture settings. As shown in Fig. 2, our network architecture consists of four stages, each incorporating a geometric affine module and several Mamba layers. We employ a simple decoder without Mamba layers for point cloud segmentation. The decoder solely conducts point cloud interpolation, concatenation with multi-stage encoder features, and channel transformation through MLP. The number of Mamba layers, serialization strategies, channel sizes, and order prompt counts in each stage of PCM-Tiny are illustrated in Tab. 1. Furthermore, we obtain PCM by increasing the number of Mamba layers and channel sizes and adopting new serialization methods accordingly.
4 Experiments
4.1 Experiment Setup
Table 2: 3D object classification in ScanObjectNN and ModelNet40. Averaged results in three random runs using $1024$ points as input without normalization and voting are reported.
@@ 4.1 Experiment Setup
PointNet [38] 68.2 63.4 89.2 86.2 3.5
PointCNN [27] 78.5 75.1 92.2 88.1 0.6
KPConv [51] - - 92.9 - 14.3
ASSANet-L  [40] - - 92.9 - 118.4
CurveNet [64] - - 93.8 - 2.0
PointMLP [35] 85.4$\pm$1.3 83.9$\pm$1.5 94.1 91.3 13.2
PointNet++ [39] 77.9 75.4 91.9 - 1.5
PointNeXt [41] 87.7$\pm$0.4 85.8$\pm$0.6 93.2$\pm$0.1 90.8$\pm$0.2 1.4

### v4

# Point Cloud Mamba: Point Cloud Learning via State Space Model

@@ Abstract
Recently, state space models have exhibited strong global modeling capabilities and linear computational complexity in contrast to transformers. This research focuses on applying such architecture to more efficiently and effectively model point cloud data globally with linear computational complexity. In particular, for the first time, we demonstrate that Mamba-based point cloud methods can outperform previous methods based on transformer or multi-layer perceptrons (MLPs). To enable Mamba to process 3-D point cloud data more effectively, we propose a novel Consistent Traverse Serialization method to convert point clouds into 1-D point sequences while ensuring that neighboring points in the sequence are also spatially adjacent. Consistent Traverse Serialization yields six variants by permuting the order of x, y, and z coordinates, and the synergistic use of these variants aids Mamba in comprehensively observing point cloud data. Furthermore, to assist Mamba in handling point sequences with different orders more effectively, we introduce point prompts to inform Mamba of the sequence’s arrangement rules. Finally, we propose positional encoding based on spatial coordinate mapping to inject positional information into point cloud sequences more effectively. Point Cloud Mamba surpasses the state-of-the-art (SOTA) point-based method PointNeXt and achieves new SOTA performance on the ScanObjectNN, ModelNet40, ShapeNetPart, and S3DIS datasets. It is worth mentioning that when using a more powerful local feature extraction module, our PCM achieves 79.6 mIoU on S3DIS, significantly surpassing the previous SOTA models, DeLA and PTv3, by 5.5 mIoU and 4.9 mIoU, respectively.
@@
Figure 1: Several pipelines of point cloud modeling. (a) denotes point-based methods with only local perception, including point-based methods, such as PointNet (Qi et al. 2017a), PointNet++ (Qi et al. 2017b), PointMLP (Ma et al. 2022), and PointNeXt (Qian et al. 2022). (b) is the transformer-based method with global perception but quadratic computational cost, including Point Transformer (Zhao et al. 2021) and Point-MAE (Pang et al. 2022). (c) represents Mamba-based methods, which offer advantages of global modeling and linear computational complexity.
Introduction
@@
Point cloud analysis (Qi et al. 2017a, b; Li et al. 2018; Zhao et al. 2021; Qian et al. 2022; Wu, Qi, and Fuxin 2019; Qian et al. 2022) has become a popular topic in 3D understanding and has drawn attention from the research community. Unlike 2D image processing, point clouds are composed of unordered and irregular point sets, making it difficult to apply the 2D image processing methods directly. Thus, recent deep-learning-based approaches (Qi et al. 2017a, b; Li et al. 2018; Wu, Qi, and Fuxin 2019; Ma et al. 2022; Qian et al. 2022; Zhao et al. 2021; Guo et al. 2021) propose using various methods, such as voxel-based and point-based, for point cloud representation. The representative works, Point-Net (Qi et al. 2017a) and Point-Net++ (Qi et al. 2017b), adopt the MLP-based design with deep hierarchical local priors, as shown in Figure 1 (a). After that, many research works (Ma et al. 2022; Boulch 2020; Shi and Rajkumar 2020; Wang et al. 2019; Zhao et al. 2021; Guo et al. 2021; Wu et al. 2022, 2024) focus on advanced local geometric modeling via convolution, graph modeling, or attention.
@@
Meanwhile, with the rapid progress of vision transformers, several works (Zhao et al. 2021; Guo et al. 2021; Wu et al. 2022, 2024) enhance the global modeling with transformer structure in point could, as shown in Figure 1 (b). In addition, transformer architectures also work effectively in mask point pre-training, 3D segmentation, and in-context learning. However, the computation and memory costs are still huge. Recently, state space models (Gu and Dao 2023; Gu, Goel, and Ré 2022) (SSMs) have been proven to model long-range dependency in sequential data. In particular, Mamba (Gu and Dao 2023) is proven effective as Transformer (Vaswani et al. 2017) for several challenging NLP tasks. After that, recent works (Zhu et al. 2024; Liu et al. 2024; Li, Singh, and Grover 2024; Behrouz and Hashemi 2024; Ma, Li, and Wang 2024; Xing et al. 2024; Yang, Xing, and Zhu 2024; Ruan and Xiang 2024; He et al. 2024) explore SSMs in various vision tasks, including image representation learning, medical segmentation, and low-level vision tasks. One concurrent work, PointMamba, uses the Mamba layer to model the global context. However, there are still significant performance gaps between PointMamba and previous point-based methods.
@@
In this work, we ask an essential question: Can we design an efficient point cloud analysis architecture using Mamba and surpass the performance of point-based and transformer-based methods? In particular, we introduce the Point Cloud Mamba (PCM), a combining local and global modeling framework that outperforms the SOTA point-based and transformer-based methods.
@@
PCM utilizes Mamba architecture to model the global features of point clouds while maintaining linear computational complexity. To process 3-D point cloud data effectively by Mamba layers, we propose a novel Consistent Traverse Serialization (CTS) method to serialize point clouds into a 1-D point sequence while ensuring that neighboring points in the sequence are also adjacent in space. Then, CTS can easily derive six variants by simply permuting the order of x, y, and z coordinates. Additionally, when these six variants of CTS are combined, Mamba layers can more effectively model point cloud features. This is because the different variants provide various perspectives of the point cloud, ensuring that spatially adjacent points are also adjacent in a serialized point sequence. To help Mamba handle specific point sequences better, we introduce order prompts to provide Mamba with the arrangement rules of the current point sequence. Finally, we propose simple spatial coordinate mapping as positional embedding for points, more suitable for irregular point cloud data than RoPE (Su et al. 2024) and learnable embedding.
@@
Thanks to the above improvements, we successfully introduced Mamba into Point Cloud analysis and obtained Point Cloud Mamba (PCM). PCM outperforms the SOTA point-based method PointNeXt on three datasets: ScanObjectNet (Uy et al. 2019), ModelNet40 (Wu et al. 2015), and ShapeNetPart (Yi et al. 2016). When enhancing the local feature extraction layers, PCM achieved 79.6 mIoU on the S3DIS dataset, significantly surpassing the previous SOTA PTv3 (Wu et al. 2024) by 4.9 mIoU.
@@
In summary, we have the following contributions: 1) We introduce Mamba into point cloud analysis and construct a combined local and global modeling framework named Point Cloud Mamba. 2) We propose consistent traverse serialization, order prompts, and positional encoding based on spatial coordinate mapping to assist Mamba in better handling point cloud data. 3) Point Cloud Mamba is the first Mamba-based method that works well in point cloud analysis. It outperforms the SOTA point-based method PointNeXt and transformer-based method PTV3 on ScanObjectNet, ModelNet40, ShapeNetPart, and S3DIS datasets.
Related Work
@@
3D Point Cloud Classification. Recent works have used deep neural networks to process 3D point clouds. In particular, representative works, PointNet (Qi et al. 2017a) and PointNet++ (Qi et al. 2017b), are the pioneering point-based approaches to handle the point clouds using MLPs directly. Meanwhile, several works explore graph-based modeling to utilize 3D geometric topology. Then, several works (Wu, Qi, and Fuxin 2019; Thomas et al. 2019; Shen et al. 2018; Li et al. 2018; Xu et al. 2021a; Komarichev, Zhong, and Hua 2019; Graham, Engelcke, and Van Der Maaten 2018; Choy, Gwak, and Savarese 2019; Zhu et al. 2021) explore the local geometric features via different kernel modeling. Moreover, several works (Liu et al. 2019; Xu et al. 2021b; Xiang et al. 2021; Ran, Liu, and Wang 2022; Chen et al. 2023b; Jiang et al. 2022; Liu, Cai, and Lee 2022; Xie et al. 2020; Zhang et al. 2021; Ma et al. 2022; Lang et al. 2019; Dai and Nießner 2018; Yan et al. 2022) explore other point cloud architecture designs, including MLPs and transformers. Several works also explore different pre-training methods (Pang et al. 2022; Yu et al. 2022; Sanghi 2020; Sauder and Sievers 2019; Wang and Solomon 2019; Hou et al. 2022; Jiang et al. 2023; Wu et al. 2023b; Zhu et al. 2023; Yang et al. 2023) or in-context abilities (Fang et al. 2023; Wang et al. 2024; Wu et al. 2023a) inspired by the NLP field. Recently, state space models (Gu, Goel, and Ré 2022; Gu and Dao 2023) have achieved significant progress. Compared with transformers, they have advantages in efficient global modeling. A concurrent work (Liang et al. 2024) explores such architecture in point clouds. However, there are still significant performance gaps compared with previous point cloud methods. Our PCM shows that Mamba architecture can achieve comparable or even better results than transformer-based models.
@@
3D Visual Transformers. With the rise of the transformer in 2D version (Dosovitskiy et al. 2021; Li et al. 2023; Carion et al. 2020), several works (Lahoud et al. 2022; Guo et al. 2021; Schult et al. 2023; Sun et al. 2023; Lai et al. 2022; Liu et al. 2023; Wang 2023) also explore transformer architectures in the point cloud. Earlier works (Guo et al. 2021; Zhao et al. 2021) have focused on the point cloud process. PCT (Guo et al. 2021) performs global attention directly to each point, following the ViT (Dosovitskiy et al. 2021). However, it has memory consumption and computational complexity issues. Point Transformer (Zhao et al. 2021) solves this issue by introducing local attention. Then, the updated versions (Wu et al. 2022, 2024) explore the different architectures to improve performance and efficiency. Inspired by these studies, our works combine local point processing and a new traverse serialization strategy, which leads to better results than direct SSM traverse.
Transformer-based methods (Zhao et al. 2021; Guo et al. 2021; Wu et al. 2022, 2024; Lai et al. 2022; Robert, Raguet, and Landrieu 2023; Park et al. 2023; Wang 2023) use Transformers to model point cloud sequences and have extensively explored a lot of aspects, such as point cloud serialization (Wang 2023; Zhao et al. 2021; Guo et al. 2021; Wu et al. 2022, 2024) and point positional embedding (Wu et al. 2022; Park et al. 2023). Although it’s possible to simply implement a Mamba-based network for point cloud analysis by replacing Transformer layers with Mamba layers, due to the differences between Transformers and Mamba, substantial exploration is still needed to find the most suitable network architecture, serialization method and positional embedding strategy for Mamba-based models. This is precisely the objective of this paper.
@@
State Space Models. Inspired by continuous state space models in control systems, recently, state space models (Gu and Dao 2023; Gu, Goel, and Ré 2022) have been proven to model long-range dependency. In particular, S4 (Gu, Goel, and Ré 2022) proposes to normalize the parameter into the diagonal structure, which results in less computation cost and memory usage. After that, Mamba (Gu and Dao 2023) presents a selection mechanism that leads to better results than transformers. Recently, several works have explored such architecture in different tasks, including image classification (Zhu et al. 2024; Liu et al. 2024; Li, Singh, and Grover 2024), graph modeling (Behrouz and Hashemi 2024), medical segmentation (Ma, Li, and Wang 2024; Xing et al. 2024; Yang, Xing, and Zhu 2024; Ruan and Xiang 2024), and low-level version tasks (He et al. 2024). As a concurrent work, we further prove the potential of SSMs in the 3D point clouds, where we can achieve even better results than previous architectures.
Method
@@
The SSM-based architecture, Mamba (Gu and Dao 2023), is attractive for point cloud representation learning due to its global modeling capability and linear computational complexity. However, Mamba is designed for the causal modeling of 1-D sequences, making it difficult to be directly applied to the modeling of non-causal 3-D point cloud data, posing many challenges to be addressed. This section explores how to effectively integrate Mamba into architectures based on local modeling to capture global features.
Preliminaries
PointMLP formulation. PointMLP (Ma et al. 2022) models the representation of point clouds using simple MLP and expands the receptive field of each point through point cloud downsampling and local feature aggregation. The process can be described by Equ. 1, 2, and 3:
@@
$GAM(\{f_{i,j}^{l}\})=\alpha\odot\frac{\{f_{i,j}^{l}\}-f_{i}^{l}}{\sigma+\delta},$ (2)
$\sigma=\sqrt{\frac{1}{k\times n\times d}\sum_{i=1}^{n}\sum_{j=1}^{k}(f_{i,j}^{l}-f_{i}^{l})^{2}},$ (3)
@@
Inspired by Equ. 5 and 6, Mamba (Gu and Dao 2023), a new SSM-based model that introduces time-varying system parameters, has been introduced. Specifically, $\Delta$, $\overline{\mathbf{A}}$, and $\overline{\mathbf{B}}$ are all functions of $x_{t}$. Mamba demonstrates long sequence modeling capabilities comparable to Transformer and achieves linear computational complexity during inference following Equ. 6. However, Equ. 6 is difficult to compute in parallel. It can be expanded and implemented using global convolution to enhance the efficiency of training Mamba on GPUs:
@@
A Naive Mamba-based Point Cloud Network
Architecture. Using Mamba (Gu and Dao 2023) to replace the MLP operator in PointMLP (Ma et al. 2022) to achieve global modeling capabilities is a natural idea. However, this still requires addressing two challenges: 1) Point clouds are 3-D data, so how can we transform them into 1-D sequences? 2) Mamba is designed for causal modeling, so how can Mamba handle non-causal point cloud data? To address these challenges, we first adopt the z-order (Morton 1966) serialization method (Wang 2023) to flatten 3-D point cloud data into 1-D sequences, allowing point cloud data to be processed by Mamba. Secondly, inspired by (Zhu et al. 2024) and (Liu et al. 2024), we use bidirectional Mamba to allow each point to obtain features from any other point. At this point, we have implemented a naive Mamba-based network for the point cloud:
$f_{i}^{l+1}=Mamba(x^{l},reverse(x^{l})),$ (8)
$x^{l}=\mathcal{S}_{z}(\mathcal{A}(GAM(\{f_{i,j}^{l}|j=1,...,k\}))),$ (9)
@@ A Naive Mamba-based Point Cloud Network
Shortcomings. The naive Mamba-based network still has some shortcomings, only achieving 84.1 OA on ScanObjectNN, significantly underperforming modern point-based methods like PointNeXt. There are some reasons: 1) Using a single serialization method to convert 3-D point cloud data into 1-D point sequences significantly loses the spatial relational information between points. Specifically, even if a point is adjacent to many others in space, it can only be adjacent to adjacent points in the point sequence. Therefore, employing multiple serialization methods in combination to alleviate this loss of spatial relational information is necessary. 2) The architecture design of the naive Mamba-based network follows the point-based method PointMLP, which may not be optimal or even reasonable for Mamba-based architecture. Exploring what benefits Mamba in modeling point clouds is important and necessary.
Point Cloud Mamba
In the last section, we obtained a naive mamba-based point cloud network; however, there is still plenty of room for optimization. Next, we will elaborate on improving this naive architecture to Point Cloud Mamba (PCM) and achieving performance beyond PointNeXt (Qian et al. 2022) and PTv3 (Wu et al. 2024).
@@
Serialization strategy. How to convert 3-D point cloud data into a 1-D sequence that Mamba can handle is crucial. We find that Mamba can better process point cloud sequences arranged according to specific rules than disordered point cloud sequences. For example, randomly flattening point cloud data into a 1-D point sequence and feeding it into Mamba for modeling will significantly underperform compared to using an ordered point sequence (84.1 vs. 86.7). Based on this, we propose the Consistent Traverse Serialization (CTS) strategy, which ensures that adjacent points in the sequence are also adjacent in spatial position. Figure 3 illustrates how CTS converts 3-D point clouds into a 1-D sequence.
@@
$Code\_func(n_{1},n_{2})=\left\{\begin{aligned} n_{2}\times N+n_{1},&n_{2}\%2=0\\ (n_{2}+1)\times N-n_{1},&n_{2}\%2\neq 0\\ \end{aligned}\right.$ (11)
@@
Sorting the point cloud according to the code allows it to be flattened into a 1-D sequence. This simple serialization strategy performs comparably to carefully designed z-order (Morton 1966) and Hilbert-order (Hilbert and Hilbert 1935) serialization strategies (Wang 2023; Wu et al. 2024). Additionally, by exchanging the order of the x, y, and z axes, six different serialization methods can be derived, which we call "xyz", "xzy", "yxz", "yzx", "zxy", and "zyx". These different serialization methods can be viewed as various point cloud observations from different spatial perspectives.
Additionally, combining multiple serialization strategies can effectively assist Mamba in better modeling point cloud features, as shown in Table 9. Specifically, we adopt different serialization strategies for the inputs of different Mamba layers, allowing Mamba to perceive the point cloud more comprehensively, thus significantly surpassing a single serialization strategy alone.
Figure 4: The order prompts. Different colors represent different serialization orders. $N_{p}$ order prompts are mapped to the same channel size as the features and then concatenated to the beginning and end of the input point sequence.
@@
As shown in Figure 4, we assigned $N_{p}$ learnable embeddings as order prompts for each serialization order. Before being processed by the Mamba layer, the point cloud sequence has the corresponding order prompts added to both the beginning and end of the sequence. Considering that the input feature dimensions of the Mamba layers at different stages may vary, we also allocate a shared Linear layer for all Mamba layers within each stage to map the order prompts to the required channel size.
Positional embedding. In sequence modeling, positional encoding is crucial and widely applied in text and image patch sequence modeling. For example, RoPE (Su et al. 2024) and learnable positional embedding are widely used in text sequence modeling and image patch sequence modeling, respectively. However, we find that these positional encoding methods are unsuitable for point clouds due to their sparsity and irregular shapes. The spatial distances between any two adjacent points in a point cloud sequence vary significantly, making it difficult to use the sequence positional information to represent the spatial gaps between adjacent points accurately. We employ a shared positional mapping function to address this challenge. It maps the point coordinates to positional embeddings with the same channel size as the features.
$Emd_{pos}=Linear(\{c_{1}^{s},c_{2}^{s},c_{3}^{s}\}).$ (13)
@@
Architecture settings. As shown in Figure 2, our network architecture consists of four stages, each incorporating a geometric affine module and several Mamba layers. We employ a simple decoder without Mamba layers for point cloud segmentation. The decoder solely conducts point cloud interpolation, concatenation with multi-stage encoder features, and channel transformation through MLP. The number of Mamba layers, serialization strategies, channel sizes, and order prompt counts in each stage of PCM-Tiny are illustrated in Table 1. Furthermore, we obtain PCM by increasing the number of Mamba layers and channel sizes and adopting new serialization methods accordingly.
Experiments
@@
PointNet (Qi et al. 2017a) 68.2 63.4 89.2 86.2 3.5
PointCNN (Li et al. 2018) 78.5 75.1 92.2 88.1 0.6
KPConv (Thomas et al. 2019) - - 92.9 - 14.3
ASSANet-L  (Qian et al. 2021) - - 92.9 - 118.4
CurveNet (Xiang et al. 2021) - - 93.8 - 2.0
PointMLP (Ma et al. 2022) 85.4$\pm$1.3 83.9$\pm$1.5 94.1 91.3 13.2
PointNet++ (Qi et al. 2017b) 77.9 75.4 91.9 - 1.5
PointNeXt (Qian et al. 2022) 87.7$\pm$0.4 85.8$\pm$0.6 93.2$\pm$0.1 90.8$\pm$0.2 1.4